# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Enhancement Protocol

### v1

@@ V-C Results
Figure 5: Boxplot depicting the distribution of absolute drops in pass@1 score between HumanEval and the newly introduced HumanEvalNext benchmark, based on 10 LLMs (Figure 6).
@@
Upon conducting the experiments outlined in subsection V-B, a key observation is an average decrease of 31.22% and a median decrease of 26.02% (both absolute percentages) in pass@111 results of HumanEval when contrasted with the newly introduced HumanEvalNext benchmark, using 10 different LLMs. Specific model outcomes are detailed in Figure 6, with an overall depiction of the performance decreases displayed in Figure 5. While there are still significant performance declines when comparing HumanEvalPlus with the original HumanEval (11.28 mean and 7.05 median), the declines are substantially larger with HumanEvalNext than with EvalPlus. Overall, the findings highlight a marked reduction in model performance when measured against HumanEvalNext as opposed to the initial HumanEval benchmark. This pattern emphasizes the enhanced difficulty and refined evaluation precision offered by HumanEvalNext, which incorporates more resilient environments featuring type annotations and clearer instructions.
@@
Figure 6: Comparison of pass@1 scores between the original HumanEval benchmark and HumanEvalNext.
VI Discussion
VI-A Implications
Our examination of the current AI4SE benchmarking environment leads us to a series of standardized insights about the domain, which we outline below.
  1. 1.
    Limited Programming Language Variety: There is a strong tendency for new benchmarks to follow the norms of the field which revolve around well-known languages like Python, Java, or C++; With this, more niche languages such as Haskell, Visual Basic, and Julia. Although projects like the MultiPL-E [6] series of automatically translated benchmarks help reduce this disparity, a lack of benchmarks focused on native low-resource languages remains.
  2. 2.
    Limited Natural Language Variety: Similar to programming languages, most benchmarks focus mainly on English, with Chinese being the next most common. Although resources like MicrosoftDocs [173] include languages like Norwegian, Danish, and Latvian, evaluation datasets still lack sufficient natural language diversity, restricting the insights into how these models perform in different languages.
  3. 3.
    Limited Task Diversity: Over the years, tasks like Text2Code have been subjected to numerous evaluation benchmarks, whereas tasks like domain-specific code generation and memorization have been largely overlooked; This, highlights an imbalance in benchmark availability.
  4. 4.
    Real-World Applicability Issues: Benchmarks traditionally centered on isolated tasks such as function-level code generation, which limited the practicality of their results in real-world scenarios. Recently, this has shifted towards more comprehensive software engineering and project-level tasks to enhance the real-world relevance of benchmark results.
@@ VI-A Implications
With tools such as BenchScout, users can select more relevant benchmarks and gain deeper insights into model performance for their specific needs. Furthermore, our results show the significant impact of integrating the BenchFrame framework and highlight the value of peer-reviewed, validated benchmarks. With many state-of-the-art models tested on similar benchmarks, both industry and academia should adapt their evaluation methods to ensure robust results. The refinement of benchmarks across various AI4SE tasks will be critical for guiding future research and ensuring that these models can perform effectively. Although this study highlights the issue of data leakage in current benchmarks, it remains true that HumanEvalNext would also be affected by this issue. Nonetheless, we argue that benchmarks should not be employed indefinitely and ought to evolve to pose increasing challenges in alignment with model improvements.
VI-B Future Work
@@
Future work on BenchFrame will focus on expanding to additional programming languages, which it already has been optimized for, yet these variants have not been produced or evaluated. Additionally, given that BenchFrame is a generalizable framework, future work would focus on applying rigorous peer review to other common benchmarks such as the MBPP benchmark family. While evaluating ten LLMs provides valuable insights, it remains unclear how larger, top-performing models behind paywalls, such as GPT and Gemini, would perform; Future research could evaluate the performance differences of such models when comparing the base and the pro version.
VI-C Threats to the Validity
@@
Internal Validity: A potential threat in the user study is selection bias. To mitigate this, we included participants from both industry and academia, ensuring a range of skills and experience. All participants received the same tool, guidance, and instructions. To reduce bias in the peer review, the reviewer was not informed of specific changes made by the first author.
External Validity: A threat here is the generalizability of the user study’s results. We mitigated this by including 22 participants, but the sample size may still limit the generalizability of the results. While we evaluated the effects of the BenchFrame framework on the performance of ten models with one of the most used benchmarks, this might not be seen as sufficient to prove the generalizability of the results; adding more models and application to more benchmarks could further confirm our results. BenchFrame may be criticized for its limited applicability due to the significant manual effort it demands compared to more automated systems like EvalPlus. Nevertheless, considering its demonstrated potential and the discrepancies in outcomes relative to automated approaches, these efforts can be deemed worthwhile.
VII Conclusion
@@
VII-A Data Avalaibilty
We publicly release the results of our literature review, user study, and 50% of the manually refined benchmark to demonstrate the quality of the data.  [^20] https://github.com/AISE-TUDelft/AI4SE-benchmarks Upon acceptance, the complete benchmark will be made available on both GitHub and HuggingFace. Because of space constraints, more detailed tables from our review are available in the appendix and also to be found in the replication package.

### v2

@@ 5.3. Results
Figure 7. Boxplot depicting the distribution of absolute drops in pass@1 score between HumanEval and the newly introduced HumanEvalNext benchmark, based on 10 LLMs (Table 22).
@@
Upon conducting the experiments outlined in Section 5.2, a key observation is an average decrease of 31.22% and a median decrease of 26.02% (both absolute percentages) in pass@$1$ results of HumanEval when contrasted with the newly introduced HumanEvalNext benchmark, using 10 different LLMs. Specific model outcomes are detailed in Table 22, with an overall depiction of the performance decreases displayed in Figure 7. While there are still significant performance declines when comparing HumanEvalPlus with the original HumanEval (11.28 mean and 7.05 median), the declines are substantially larger with HumanEvalNext than with EvalPlus. Overall, the findings highlight a marked reduction in model performance when measured against HumanEvalNext as opposed to the initial HumanEval benchmark. This pattern emphasizes the enhanced difficulty and refined evaluation precision offered by HumanEvalNext, which incorporates more resilient environments featuring type annotations and clearer instructions.
@@
Table 22. Performance Comparison on HumanEval Benchmarks (pass@1)
Model HumanEval HumanEvalPlus ($\Delta$) HumanEvalNext ($\Delta$)
stable-code-3b 30.72 25.60 (5.12-5.12) 1.83 (28.89-28.89)
CodeLlama-13b-instruct-hf 50.60 34.10 (16.50-16.50) 29.88 (20.72-20.72)
codegemma-7b-it 60.40 51.80 (8.60-8.60) 41.46 (18.94-18.94)
starchat2-15b-v0.1 73.80 71.30 (2.50-2.50) 43.29 (30.51-30.51)
Artigenz-Coder-DS-6.7B 75.60 72.60 (3.00-3.00) 53.66 (21.94-21.94)
Magicoder-S-DS-6.7B 76.80 71.30 (5.50-5.50) 53.66 (23.14-23.14)
starcoder2-15b-instruct 77.40 60.00 (17.40-17.40) 43.29 (34.11-34.11)
deepseek-coder-6.7b-instruct 80.22 71.30 (8.92-8.92) 58.54 (21.68-21.68)
CodeQwen1.5-7B 87.20 45.70 (41.50-41.50) 10.98 (76.22-76.22)
Nxcode-CQ-7B-orpo 87.23 83.50 (3.73-3.73) 51.22 (36.01-36.01)
6. Discussion
In this section, we discuss the broader implications of our study; here, we attempt to tie our findins back to the RQs that guided our study. We, specifically, touch on what our results mean for researchers navigating the field, address the practical salability of our proposed framework, and outline threats to validity and opportunities for future work.
6.1. Implications
6.1.1. Finding the Right Benchmark
In our analysis, and specifically RQ1 we find that AI4SE benchmarking is highly fragmented, with many benchmarks suffering from poor maintenance and lack of discoverability. With this finding, we underscore the fact a major challenge in AI4SE is the difficulty of selecting a contextually relvant benchmark. BenchScout was our direct answer to RQ2. With BenchScout, users can select more relevant benchmarks and gain deeper insights into model performance for their specific needs.
6.1.2. Benchmarking the Benchmark
@@
Our answer to RQ3, BenchFrame, can have an impact on how we measure progress in AI4SE. Our results show the significant impact of integrating the BenchFrame framework and highlight the value of peer-reviewed, validated benchmarks. With many state-of-the-art models tested on similar benchmarks, both industry and academia should adapt their evaluation methods to ensure robust results. The refinement of benchmarks across various AI4SE tasks will be critical for guiding future research and ensuring that these models can perform effectively. Although this study, specifically in RQ1, highlights the issue of data leakage in current benchmarks, it remains true that HumanEvalNext would also be affected by this issue. However, we argue that benchmarks should not be used indefinitely and should evolve to pose increasing challenges in alignment with model improvements.
6.2. A Case Study on the agentification of BenchFrame
A potential criticism of the framework proposed in RQ3 is the significant manual effort it demands compared to a more automated system. To address these concerns about the generalizability and manual labor intensity of BenchFrame, we developed an agentic pipeline designed to replicate and automate the benchmark improvement process with minimal human intervention. Our baseline for this effort was the HumanEvalNext dataset, which had previously been improved by manual human supervision. We aimed to determine whether an agentic approach could yield comparable improvements without the extensive human labor (i.e., to show noninfiriority).
The agent pipeline was structured into three distinct phases, which mirrors the manual process described in Section 5.1.2. In the first phase, the agent improved the problem description, which includes the function signature and the docstring. Using this refined description, the agent then generated an improved canonical solution. In the final phase, the agent created and set up the test cases to validate the solution. This process was applied across all the problems in the dataset, and our hypothesis was that this automated approach would be noninferior to the human-improved dataset.
To validate this, we conducted a paired evaluation comparing the human- and agent-improved versions of each benchmark problem in the HumanEvalNext dataset, specifically HumanEvalNext versus HumanEvalNext-Agentic. Two independent reviewers rated each pair using a five-point ordinal scale ranging from -2 (strongly preferring the human version) to +2 (strongly preferring the agentic version), with 0 indicating no preference. We tested whether the agentic pipeline was noninferior to the human-improved process, using a conservative noninferiority margin of $\delta$ = -0.5. That is, we assessed whether the average rating for the agentic outputs was not more than 0.5 points worse than the human outputs.
For each reviewer, as well as for the combined ratings (averaged across reviewers), we performed both a one-sided one-sample t-test and a Wilcoxon signed-rank test to assess noninferiority. The results were consistent and conclusive: the mean ratings were 0.16 (SD = 0.67) for Reviewer 1, 0.53 (SD = 0.51) for Reviewer 2, and 0.35 (SD = 0.49) for the combined average. The corresponding t-statistics were 12.68, 25.75, and 22.10, with one-sided p-values all equal to 1.000 (indicating extremely strong evidence in favor of noninferiority given the direction of the test). The non-parametric Wilcoxon signed-rank tests further confirmed these results, with statistics of 12,155.5 (p = $3.46\times 10^{-20}$), 13,198.0 (p = $7.20\times 10^{-28}$), and 12,698.0 (p = $5.94\times 10^{-27}$) respectively.
These results indicate that the agentic pipeline is noninferior to the human-improved process—indeed, with mean and median ratings above zero, the agentic output was often preferred. This supports the feasibility of using an agentic approach to automate benchmark refinement while significantly reducing manual effort.
While noninferior, we noted several standardized observations, in terms of pitfalls, during this evaluation, such as (1) instances where the agent’s docstring occasionally revealed aspects of the solution, or (2) where canonical solutions made unintended assumptions, or (3) where the test cases did not adhere to the required, single-line assert, format. These observations were later accounted for in later stages of the feasibility study to make adjustments in the pipeline, and to ensure that it remained flexible and effective.
Initialize attempt = 1Text Improvement (a)Code Improvement (b)Test Improvement (c)All tests pass?Validation (d)attempt $\leq$ 3attempt = attempt + 1Minimized passing test cases (e)Mark task as failedyesnoyesnoyesno
Figure 8. A flow-chart indicating the pipeline used to improve MBPP.
For the sake of completeness, we present the refined pipeline used in the process of benchmark refinement. The pipeline begins by initializing an attempt counter, which tracks the number of iterations taken to complete a given benchmark improvement. The process then proceeds through three core phases: (a) Text Improvement, in which the agent refines the problem description and adds a type-annotated function signature; (b) Code Improvement, where the canonical implementation is modified according to Python best practices and based on the improved task description; and (c) Test Improvement, which involves generating assert-based test cases that include edge cases, boundary conditions, and common error scenarios. These three components are then assembled into a complete Python program and passed to the Validation phase (d), which runs the implementation against the test suite in a sandboxed subprocess.
If all tests pass on the first try, the process terminates successfully. Otherwise, the pipeline evaluates whether the number of attempts is below a fixed threshold (three, in our case). If so, the pipeline either increments the attempt counter and retries the full process or, alternatively, in the case that more tests are passing than were present in the original test suite, the pipeline identifies a minimal set of passing tests (e) to be used as the new test suite. This loop ensures that even partially successful test runs can be used to improve the benchmark iteratively. After three failed attempts, the task is marked as unsuccessful.
For our experiments, we used the OpenAI API  [^22] https://platform.openai.com, specifically, the o3-mini-2025-01-31 as the underlying model for the agent. We evaluated this new dataset on the same set of models, and the results (as presented in Table 23) indicated comparable (and even better) performance patterns. These findings provide adequate justification for how an agentic pipeline can effectively generalize and provide a semi-automated solution for benchmark improvement.
Table 23. Pass@1 results for HumanEval, HumanEvalNext, and HumanEvalNext-Agentic.
Model HumanEval HumanEvalNext-Agentic ($\Delta$) HumanEvalNext ($\Delta$)
stable-code-3b 30.72 2.44(28.28-28.28) 1.83 (28.89-28.89)
CodeLlama-13b-instruct-hf 50.60 18.90(31.70-31.70) 29.88 (20.72-20.72)
codegemma-7b-it 60.40 22.56 (37.84-37.84) 41.46 (18.94-18.94)
starchat2-15b-v0.1 73.80 26.83 (46.97-46.97) 43.29 (30.51-30.51)
Artigenz-Coder-DS-6.7B 75.60 31.70 (43.90-43.90) 53.66 (21.94-21.94)
Magicoder-S-DS-6.7B 76.80 31.70 (45.10-45.10) 53.66 (23.14-23.14)
starcoder2-15b-instruct 77.40 14.02 (63.38-63.38) 43.29 (34.11-34.11)
deepseek-coder-6.7b-instruct 80.22 35.37 (44.85-44.85) 58.54 (21.68-21.68)
CodeQwen1.5-7B 87.20 1.22 (86.00-86.00) 10.98 (76.22-76.22)
Nxcode-CQ-7B-orpo 87.23 31.70 (55.53-55.53) 51.22 (36.01-36.01)
The results presented indicate, that while a certain level of human intervention, specifically in the review phase, is still required to ensure quality of the improvements, it is but a fraction of the total time required to manually improve the benchmark. From a cost perspective, we can also report that the total incurred cost from calling the apis for the models was \$5.28232323At the time of running these experiments, on the 31st of July 2025. This is but a small portion of the labor costs required to improve the benchmark manually. The introduction of this agentic layer significantly reduced the manual effort required. We claim that, with this approach, one can extend it to other datasets with similar success.
6.3. Evaluating the Generalizability of BenchFrame
An additional critique that can be raised regarding our approach is its generalizability. To tackle this issue, we adopted a strategy akin to HumanEval. Using our systematic observations of MBPP from the review, we apply BenchFrame to 100 problems selected from a pool of 500 MBPP test samples. The rationale behind this selection is twofold: firstly, it aims to assess a new task, specifically program synthesis, and secondly, it considers the benchmark’s popularity. During this process, we subject the function descriptions, implementations (including variable typing), and significantly extend test cases; more specifically, we go from a fixed 3 tests per problem to an average of 12.35, with a median of 11.5. We finally go through several rounds of peer review to ensure the quality of the improvements. We refer to the resulting benchmark as MBPPNext. The evaluation results of this dataset in comparison to the baseline are detailed in Table 24. We observe a comparable decline in performance similar to that noted for HumanEval, specifically, an average decrease of 13.4 percentage points across the methods evaluated. These findings further underscore the necessity for thorough peer review and verification of benchmarks, as discussed in previous sections of this paper.
Table 24. Pass@1 Results for MBPP and MBPPNext. Based on https://github.com/google-research/google-research/blob/master/mbpp/mbpp.jsonl accessed on the 14th of July 2025
Model MBPP MBPPNext ($\Delta$)
stable-code-3b 34.0 28.0 (-6%)
CodeLlama-13b-instruct-hf 40.0 31.0 (-9%)
codegemma-7b-it 43.0 33.0 (-10%)
starchat2-15b-v0.1 59.0 42.0 (-17%)
Artigenz-Coder-DS-6.7B 60.0 47.0 (-13%)
Magicoder-S-DS-6.7B 63.0 45.0 (-18%)
starcoder2-15b-instruct 44.0 36.0 (-8%)
deepseek-coder-6.7b-instruct 68.0 50.0 (-18%)
CodeQwen1.5-7B 53.0 40.0 (-13%)
Nxcode-CQ-7B-orpo 74.0 52.0 (-22%)
6.4. Future Work
@@
Future work on BenchFrame will focus on expanding to additional programming languages, which it already has been optimized for, yet these variants have not been produced or evaluated. Although evaluating ten LLMs provides valuable insight, it remains unclear how larger, top-performing models behind paywalls, such as GPT and Gemini, would perform; Future research could evaluate the performance differences of such models when comparing the base and the pro version. Furthermore, future research should focus on applying the underlying ideas of BenchFrame to mutli-file and project-level benchmarks like Defects4J.
6.5. Threats to the Validity
@@
Internal Validity: A potential threat in the user study is selection bias. To mitigate this, we included participants from both industry and academia, ensuring a range of skills and experience. To reduce the real-world impact arising from any study biases, we have created a continuous feedback mechanism in BenchScout to receive continuous feedback from users to be able to meet their needs. All participants received the same tool, guidance, and instructions. To reduce bias in the peer review or HumanEvalNext, the reviewer was not informed of specific changes made by the first author.
External Validity: A threat here is the generalizability of the user study’s results. We mitigated this by including 22 participants, but the sample size may still limit the generalizability of the results. Although we evaluated the effects of the BenchFrame framework on the performance of ten models with one of the most widely used benchmarks, this could not be considered sufficient to prove the generalizability of the results; adding more models and application to more benchmarks could further confirm our results. BenchFrame may be criticized for its limited applicability due to the significant manual effort it demands compared to more automated systems like EvalPlus. We demonstrate that the process of improving benchmarks such as HumanEval and MBPP can be (semi-)automated through the use of agentic pipelines.
7. Related Work
The rapid integration of LLMs into software engineering has led to a corresponding proliferation of benchmarks designed to evaluate them. This has spurred research to systematically map this new terrain, assess the quality of evaluation resources, and build tools for navigating them. Our work is situated within these three emerging areas.
AI4SE Surveys and Taxonomies
Initial research in this area included broad surveys of LLMs in software engineering, which established evaluation as a central research challenge (Fan et al., 2023b; Hou et al., 2024; Zakeri-Nasrabadi et al., 2023; Zhang et al., 2023c). More recently, meta-analyses have focused on the benchmarks themselves. For instance, Wang et al. created a taxonomy based on the Software Development Life Cycle, revealing a research gap with a strong focus on code implementation and a negligence towards design and requirements engineering (Wang et al., 2025a). Our work contributes to this area by providing a comprehensive review of 273 benchmarks from 247 studies (RQ1). We go beyond categorization by analyzing each benchmark across a 14-point metadata schema to identify systemic limitations, such as poor maintenance, language specificity, and a lack of peer review.
Tools for Benchmark Discovery and Navigation
While platforms like Hugging Face and the now-discontinued Papers With Code serve as valuable repositories for hosting datasets (Hugging Face, 2016), they often lack the specialized search and visualization capabilities needed by AI4SE researchers. BenchScout addresses this gap (RQ2) by providing an extensible semantic search tool specifically designed for the AI4SE community. Unlike general-purpose repositories, it combines semantic embeddings with structured metadata filters and an interactive visualization of the benchmark landscape to improve the discoverability and selection of relevant evaluation tools.
Methodologies for Improving Benchmark Quality
There is a growing agreement that many static benchmarks are prone to saturation and data contamination, which can compromise the validity of evaluation results (Jain et al., 2024). In response, two movements have arisen. Efforts like LiveCodeBench fall under the ”build anew” category; these strive to develop dynamic, ”live” benchmarks that continuously gather new challenges from real-world environments to avoid overfitting.
Our method, BenchFrame, belongs to the ”repair and refine” category and focuses on improving the quality of existing benchmarks (RQ3). A method within this category is DyCodeEval, which uses seed contexts to frame problems in specific scenarios (e.g., banking, healthcare, education). This approach is designed to produce more reliable results by minimizing data contamination and memorization effects (Chen et al., 2025). Our approach sets itself apart by incorporating a peer-review-oriented process that includes standardized observations, precise modifications, and independent validation. Like DyCodeEval, we show its practical scalability through an agentic pipeline that automates the improvement process and tackles the main challenge, namely, the manual efforts required in BenchFrame.
8. Conclusion
@@
8.1. Data Availability
We publicly release the results of our literature review, user study, and 50% of the manually refined benchmark.  [^24] https://github.com/AISE-TUDelft/AI4SE-benchmarks Upon acceptance, the complete benchmark will be made available on both GitHub and HuggingFace.